Hello. I am trying to figure out these 4 problems...
2x+5=x^2
x^2+3x=5
f(x) = 4(x-3)^2+5
Find quadratic function (0,0), (3,0), and (5,2)
thank you.

Answer by lynnlo(4176)

(1)y=x^2/5-3x/5
(2)x=1+√6,1 or -√6
(3)x=-3+√29/2,-3-√29/2
(4)vertex=(3,5) focus=(3,81/16)
directrix=y=79/16
axis of symmetry=x=3